# Dynamics-aware Diffusion Models for Planning and Control

### v2

@@ Abstract
This paper addresses the problem of generating dynamically admissible trajectories for control tasks using diffusion models, particularly in scenarios where the environment is complex and system dynamics are crucial for practical application. We propose a novel framework that integrates system dynamics directly into the diffusion model’s denoising process through a sequential prediction and projection mechanism. This mechanism, aligned with the diffusion model’s noising schedule, ensures generated trajectories are both consistent with expert demonstrations and adhere to underlying physical constraints. Notably, our approach can generate maximum likelihood trajectories and accurately recover trajectories generated by linear feedback controllers, even when explicit dynamics knowledge is unavailable. We validate the effectiveness of our method through experiments on standard control tasks and a complex non-convex optimal control problem involving waypoint tracking and collision avoidance, demonstrating its potential for efficient trajectory generation in practical applications.
@@ II Problem formulation and preliminaries
[…]
@@ II-A Denoising Diffusion Probabilistic Models (DDPMs)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionClean trajectory $\tau_{0}$oised sample $\tau_{L}\sim\mathcal{N}(0,I)$Generated trajectory $\tau^{\prime}_{0}$$\tau_{0}(0)$$\tau_{0}(T)$$\tau_{i}(0)$$\tau_{i}(T)$$\tau_{i}(0)$$\tau_{i}(T)$$\tau^{\prime}_{0}(0)$$\tau^{\prime}_{0}(T)$Forward diffusionReverse denoising
@@ IV Experimental Results
We evaluate our proposed method on a test bed of different control problems. We compare our methods against a vanilla diffusion model with no dynamics awareness. Our experiments were conducted on an Intel i9-9900 machine with 128GB RAM and an Nvidia Quadro RTX 4000 GPU.
@@ IV-A LQR with 4 dimensional system
[…]
@@
Refer to caption(a)time (t)Avg. $\|x(t)-x_{\text{LQR}}(t)\|^{2}$efer to captiontime (t)(b)Avg. $\|u(t)-u_{\text{LQR}}(t)\|^{2}$
@@ IV-B Waypoint tracking and collision avoidance
Refer to caption(a)x positiony positionRefer to caption(b)x positionRefer to caption(c)time (t)Avg. $\|x(t)-x_{\text{LQR}}(t)\|^{2}$efer to captiontime (t)(d)Avg. $\|u(t)-u_{\text{LQR}}(t)\|^{2}$

### v3

@@ Abstract
This paper addresses the problem of generating dynamically admissible trajectories for control tasks using diffusion models, particularly in scenarios where the environment is complex and system dynamics are crucial for practical application. We propose a novel framework that integrates system dynamics directly into the diffusion model’s denoising process through a sequential prediction and projection mechanism. This mechanism, aligned with the diffusion model’s noising schedule, ensures generated trajectories are both consistent with expert demonstrations and adhere to underlying physical constraints. Notably, our approach can generate maximum likelihood trajectories and accurately recover trajectories generated by linear feedback controllers, even when explicit dynamics knowledge is unavailable. We validate the effectiveness of our method through experiments on standard control tasks and a complex non-convex optimal control problem involving waypoint tracking and collision avoidance, demonstrating its potential for efficient trajectory generation in practical applications. Our code repository is available at www.github.com/darshangm/dynamics-aware-diffusion.
@@ II Problem formulation and preliminaries
Here, $x(0:T)=[x(0)^{\top}\ x(1)^{\top}\ \dots\ x(1)^{\top}]^{\top}$ and $u(0:T-1)=[u(0)^{\top}\ u(1)^{\top}\ \dots\ u(T-1)^{\top}]^{\top}$ denote the concatenated state and control input sequences, respectively. The environmental variable $\mathcal{E}$ encapsulates task-specific information. For instance, $\mathcal{E}$ can encode the reward function parameters, obstacle configurations, or reference trajectories. Note that the initial state $x_{0}$ is fixed. We seek to determine the optimal state and control sequence (x(0:T),u(0:T1))(x(0:T),u(0:T-1)) that maximizes the expected reward $\mathcal{R}$ while adhering to the system dynamics. This problem is challenging due to potential non-convexities and randomness in the reward function $\mathcal{R}$ introduced by the environmental variable $\mathcal{E}$. Traditional control methods relying on convexity and Gaussian assumptions may be inadequate in such scenarios.
@@ II-A Denoising Diffusion Probabilistic Models (DDPMs)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionClean trajectory $\tau_{0}$Noised sample $\tau_{L}\sim\mathcal{N}(0,I)$Generated trajectory $\tau^{\prime}_{0}$$\tau_{0}(0)$$\tau_{0}(T)$$\tau_{i}(0)$$\tau_{i}(T)$$\tau_{i}(0)$$\tau_{i}(T)$$\tau^{\prime}_{0}(0)$$\tau^{\prime}_{0}(T)$Forward diffusionReverse denoising
@@ IV Experimental Results
We evaluate our proposed method on a test bed of different control problems. We compare our methods against a vanilla diffusion model with no dynamics awareness. Our experiments were conducted on an Intel i9-9900 machine with 128GB RAM and an Nvidia Quadro RTX 4000 GPU. Our code repository is available at www.github.com/darshangm/dynamics-aware-diffusion.
@@ IV-A LQR with 4 dimensional system
which corresponds to the dynamics of a double integrator discretized with a time step of 0.1ss. The LQR problem is,
@@
Refer to caption(a)time (t)Avg. $\|x(t)-x_{\text{LQR}}(t)\|^{2}$Refer to captiontime (t)(b)Avg. $\|u(t)-u_{\text{LQR}}(t)\|^{2}$
@@ IV-B Waypoint tracking and collision avoidance
Refer to caption(a)x positiony positionRefer to caption(b)x positionRefer to caption(c)time (t)Avg. $\|x(t)-x_{\text{LQR}}(t)\|^{2}$Refer to captiontime (t)(d)Avg. $\|u(t)-u_{\text{LQR}}(t)\|^{2}$